Crane Company issues a 8%, 6-year mortgage note on January 1, 2017, to obtain financing for new equipment. Land is used as collateral for the note. The terms provide for semiannual installment payments of $45,400.
What are the cash proceeds received from the issuance of the note?

Crane Company should recieve?
Answer #1

Answer:

Semiannual payments = $45,400

Interest rate = 8%

Semiannual interest rate = 8%/2 = 4%

Maturity = 6 * 2 = 12 semiannual periods

Cash proceeds will be equal to present value above annuity.

= PV (rate, nper, pmt, fv, type)

= PV(4%, 12, -45400, 0, 0)

= $426,082.35

Cash proceeds received from the issuance of the note = $426,082.35